# You Only Forward Once: An Efficient Compositional Judging Paradigm

### v1

@@ 1 Introduction
Large language models (LLMs) [31, 38, 1, 3, 19, 10] have achieved remarkable progress, and advances in multimodal learning now enable them to process diverse modalities. Multimodal Large Language Models (MLLMs) [5, 34, 4, 21, 20] have recently demonstrated significant promise in information matching due to their ability to jointly process images, videos, and natural language queries, making them far more applicable in real-world scenarios than traditional unimodal retrievers. In particular, state-of-the-art MLLM-based rerankers harness powerful vision encoders and causal attention mechanisms to effectively fuse visual and textual information, usually ultimately regressing a single relevance score, and achieves strong performance in cross-modal matching tasks such as image search and e-commerce recommendation. However, despite these advances, outputting only a scalar relevance score leads to substantial information loss, as fine-grained semantic distinctions in complex user queries—such as compositional intent, attribute-level preferences, or contextual references—are inevitably collapsed into a single value. As an example, Fig. 1 illustrates a recommendation failure of the state-of-the-art (SOTA) multimodal reranker, Jina-Reranker-M0 [2], which prioritizes superficial semantic overlap between the user’s query and the pink dress, resulting in an inappropriately high ranking for this item. Therefore, achieving accurate and fine-grained understanding of intricate multimodal queries has become an urgent challenge that must be addressed to realize the full potential of next-generation information matching systems.
@@
Recent studies on matching can be grouped into four categories. 1) Representation-based architectures, as illustrated in Fig. 2(a), independently embed the query and the document and then compute their similarity [28, 15, 8, 30, 26, 33, 6]. However, fixed-size embeddings limit these models’ capacity, making it difficult to capture all relevant information and to meet fine-grained requirements [35]. 2) Interaction-based models have emerged to mitigate these limitations: they first compute low-level interaction representations between the query and document, and then a neural network analyzes these patterns to output a scalar relevance score for the query–document pair [13, 7, 25, 9]. These approaches achieve higher ranking quality than representation-based methods but are computationally more expensive and remain constrained by representational dimensionality. 3) MLLM/LLM-based architectures, adapted to output a single relevance score improve matching quality by leveraging the deep semantic understanding and long-context capacity of large language models [24]. However, this design is misaligned with the native generative objective of LLMs and undermines fine-grained input understanding. 4) LLMs and MLLMs can serve as native judges in a paradigm termed generative matching [40, 39], which reframes the task as autoregressive generation aligned with their generative nature. Nevertheless, these methods incur high computational overhead due to token-by-token decoding, making them difficult to deploy in real-world applications.
@@
To evaluate the efficacy of our method, we derive a training set from SA-1B [16] and train the model on it to improve generalizability. We then construct a test set from LRVS-Fashion [17] to assess recommendation performance in a different domain. Extensive experiments demonstrate that our method accurately produces judgments conditioned on the input and template, and that later judgments can condition on earlier ones, enabling dependency-aware analysis. Furthermore, YOFO also benefits from post-hoc Chain-of-Thought (CoT), which encourages the model to reason implicitly before judging.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) [5, 34, 4, 21, 20, 12, 23, 22, 36] have significantly extended the capabilities of Large Language Models (LLMs) [31, 38, 1, 3, 19, 10, 18] by enabling them to perceive, reason over, and generate responses grounded in multimodal inputs. However, the dense visual token sequences required for such alignment incur prohibitive computational costs. Recent approaches address this by either compressing tokens while preserving fidelity or rethinking how visual information is integrated. For instance, CrossLMM [37] employs dual cross-attention mechanisms to drastically reduce token count with minimal performance degradation. Similarly, MM-GEM [22] uses a PoolAggregator to support both generative and embedding tasks. These advances demonstrate that strategic token reduction can maintain multimodal performance at lower cost.
@@ 2.2 Rerankers
Reranking serves as a critical refinement stage in modern retrieval systems, aiming to re-order an initial candidate set by modeling fine-grained query–document interactions. Early neural rerankers employed representation-based architectures, calculating embeddings independently and then similarity, such as DSSM [15, 8], Convolutional DSSM [30]) and Long Short-Term Memory (LSTM) variants [11, 26, 33, 6]. But this kind of models are limited by the fixed-size embedding, and thus were soon superseded by interaction-based cross-encoders, most notably BERT-based models, that concatenate query and document as input and predict relevance via a classification or regression head over the [CLS] token. [25] first introduced this method and proved it effective. These models achieve strong performance but require pairwise inference, limiting scalability, and is constrained by fixed representational dimention and content length.   With the rise of LLMs, researchers have repurposed powerful encoder-decoder (e.g., T5 [29]) or decoder-only (e.g., LLaMA [32]) backbones for reranking. A dominant paradigm [41, 24] fine-tunes LLMs to regress continuous relevance scores or classify relevance levels, leveraging their rich contextual understanding and long-context capabilities for improved semantic matching. Jina-Reranker-M0 also follows this paradigm but leverages Qwen2-VL [34], enabling support for multimodal inputs.   More recently, generative reranking has emerged as a paradigm shift: instead of scoring candidates individually, the model directly generates the ranked sequence—such as a list of document indices or IDs—conditioned on the full candidate set. Methods like [40, 39] process all candidates jointly and autoregressively output the optimal ranking. This generative formulation naturally integrates reranking into end-to-end AI systems and shows particular promise in multimodal and agent-driven retrieval scenarios.
@@ 3.1 Problem formulation
As illustrated in Sec. 1, models akin to Jina-Reranker-M0 [2] struggle to perform accurate cross-modal retrieval. One might attribute this to visual information degradation caused by the image encoder and the connector projection, which impairs fine-grained alignment. However, we empirically find that these models also fail to retrieve text accurately from natural-language queries. As shown in Tab. 1(a), Jina-Reranker-M0 assigns a higher score to winter trousers simply because they are gray, failing to recognize that the query explicitly requires summer-appropriate clothing, while “gray” is only a conditional preference. After further simplifying the prompt, Tab. 1(b) shows that Jina-Reranker-M0 fails to handle negation reliably, appearing to measure lexical overlap rather than semantic similarity on slightly more complex inputs. Verifying each requirement in the query individually would likely yield better results. To this end, we formally define the compositional judging problem.
@@ 4 Experiments
In this section, we present extensive experiments and ablation studies to validate our approach. We first compare YOFO with Jina-Reranker-M0 [2], a representative state-of-the-art (SOTA) multimodal reranker, on the test set derived from LAION-RVS-Fashion [17]. We then assess YOFO’s dependency-aware judgement, the effectiveness of post-hoc CoT, and the contribution of individual architectural components.
@@ 4.1 Impementation Details
We build on Qwen2-VL-2B-Instruct [34] and Qwen3-VL-2B-Instruct [27], freezing the vision encoder during training. We fine-tune the LLM using LoRA [14]. To conveniently locate the unknown token, we add four special tokens to the Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct processor: <|auth_start|>, <|auth_end|>, <|reason_start|>, and <|reason_end|>. The first two mark the start and end of the answer span, and the last two mark the boundaries of the reason span. Consequently, these special tokens enable straightforward label masking, allowing supervision to be applied exclusively to answers and reasons.
@@ 4.2 Data Construction
Training set. To construct the training dataset, we use the SA-1B dataset [16], introduced alongside the Segment Anything Model. It contains approximately 11M diverse, high-resolution images sourced from a privacy-preserving data engine and annotated with 1.1B high-quality segmentation masks. Since we do not require the masks, as shown in Fig. 4(a), we use only the images, prompting an MLLM to analyze them, propose properties that hold or do not hold for the corresponding images, and provide reasons. The detailed prompt is shown in Fig. 5(a). We randomly sample 1.2M images to ensure diversity and generalizability. For each image $\mathcal{I}_{i}$, the final sample is represented as $(\mathcal{I}_{i},(\mathcal{P}_{ij},\mathcal{A}_{ij},\mathcal{R}_{ij})_{j=1}^{N})$, denoting the image, the proposed properties, the binary answers(“yes” or “no”), and the reasons. To maintain a balanced and randomized mix of satisfied and unsatisfied properties, we prompt the MLLM to produce approximately equal numbers of each and shuffle them during training. Additionally, although we instruct the model to generate exactly ten properties per sample in the training set, the number of properties per training sample varies, typically remaining close to ten.. To evaluate the model and mitigate overfitting, we split the dataset at a ratio of $6:1$ into training and validation sets.
Test set. We use the LRVS-Fashion dataset [17], a large-scale public benchmark for referred visual search (RVS) in the fashion domain, to construct a test set. It comprises 272K fashion products and 842K images extracted from real-world catalogs. To evaluate generalization in Image-Text reranking, as depicted as Fig. 4(b), we construct image pairs and queries; the model then judges which of the two images better matches the corresponding query. Specifically, we randomly sample image pairs $(\mathcal{I}_{i1},\mathcal{I}_{i2})$ from the same category and, using an MLLM, generate a natural-language query $\mathcal{Q}_{i}$ that mimics how a customer would express their preferences. The MLLM then produces the answer $\mathcal{A}_{i}$ indicating whether the first image $\mathcal{I}_{i1}$ better matches $\mathcal{Q}_{i}$, the set of query requirements $(\mathcal{R}_{ij})_{j=1}^{M_{i}}$ (where $M_{i}$ is the number of requirements), and a valid python expression $\mathcal{E}_{i}$ used to compute the final recommendation score based on our model’s predictions. The detailed prompt is demonstrated in Fig. 5(b).
@@ 4.4 Comparison on LAION-RVS-Fashion
In this section, we present an analysis of YOFO against state-of-the-art multimodal rerankers [2].
@@
YOFO also delivers substantially higher throughput than generative rerankers. This efficiency derives from YOFO’s single-forward-pass design: unlike generative or multi-stage rerankers [24, 25] that must process each requirement sequentially or in multiple rounds, YOFO produces all judgments in a single forward pass. Consequently, its inference latency increases only mildly with the number of requirements compared with generative or multi-stage rerankers, making it especially well suited to real-time, large-scale recommendation scenarios where both responsiveness and expressiveness are critical.
@@ 4.6 Ablation Study
Effectiveness of YOFO training. We fine-tune Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct models on the training set using LoRA [14] for one epoch and evaluate it on the validation set. As shown in Tab. 4, after training, Property-wise Accuracy improves by over 20%, reaching above 90%, while Sample-wise Accuracy increases by approximately 40%. These ablation results demonstrate the effectiveness of our training paradigm. Moreover, the YOFO model trained with Qwen3-VL-2B-Instruct performs slightly better than its counterpart trained with Qwen2-VL-2B-Instruct. The Sample-wise Accuracy remains moderate because the training and validation sets contain, on average, nearly ten properties per sample, making it challenging to judge all properties correctly.

### v2

@@ 1 Introduction
Large language models (LLMs) [team2024qwen2, yang2025qwen3, ahmed2025qwen, bai2023qwen, liu2024deepseek, guo2025deepseek] have achieved remarkable progress, and advances in multimodal learning now enable them to process diverse modalities. Multimodal Large Language Models (MLLMs) [bai2025qwen2, wang2024qwen2, bai2023qwenvl, lu2025ovis2, lu2024ovis] have recently demonstrated significant promise in information matching due to their ability to jointly process images, videos, and natural language queries, making them far more applicable in real-world scenarios than traditional unimodal retrievers. In particular, state-of-the-art MLLM-based rerankers harness powerful vision encoders and causal attention mechanisms to effectively fuse visual and textual information, usually ultimately regressing a single relevance score, and achieves strong performance in cross-modal matching tasks such as image search and e-commerce recommendation. However, despite these advances, outputting only a scalar relevance score leads to substantial information loss, as fine-grained semantic distinctions in complex user queries—such as compositional intent, attribute-level preferences, or contextual references—are inevitably collapsed into a single value. As an example, Fig. 1 illustrates a recommendation failure of the state-of-the-art (SOTA) multimodal reranker, Jina-Reranker-M0 [jina2025blog], which prioritizes superficial semantic overlap between the user’s query and the pink dress, resulting in an inappropriately high ranking for this item. Therefore, achieving accurate and fine-grained understanding of intricate multimodal queries has become an urgent challenge that must be addressed to realize the full potential of next-generation information matching systems.
@@
Recent studies on matching can be grouped into four categories. 1) Representation-based architectures, as illustrated in Fig. 2(a), independently embed the query and the document and then compute their similarity [radford2021learning, huang2013learning, gao2014modeling, shen2014learning, palangi2016deep, wan2016deep, cohen2016end]. However, fixed-size embeddings limit these models’ capacity, making it difficult to capture all relevant information and to meet fine-grained requirements [xu2025survey]. 2) Interaction-based models have emerged to mitigate these limitations: they first compute low-level interaction representations between the query and document, and then a neural network analyzes these patterns to output a scalar relevance score for the query–document pair [hu2014convolutional, dai2019deeper, nogueira2019multi, gao2022long]. These approaches achieve higher ranking quality than representation-based methods but are computationally more expensive and remain constrained by representational dimensionality. 3) MLLM/LLM-based architectures, adapted to output a single relevance score improve matching quality by leveraging the deep semantic understanding and long-context capacity of large language models [ma2024fine]. However, this design is misaligned with the native generative objective of LLMs and undermines fine-grained input understanding. 4) LLMs and MLLMs can serve as native judges in a paradigm termed generative matching [yoon2024listt5, yang2025rank], which reframes the task as autoregressive generation aligned with their generative nature. Nevertheless, these methods incur high computational overhead due to token-by-token decoding, making them difficult to deploy in real-world applications.
@@
To evaluate the efficacy of our method, we derive a training set from SA-1B [kirillov2023segment] and train the model on it to improve generalizability. We then construct a test set from LRVS-Fashion [lepage2023lrvs] to assess recommendation performance in a different domain. Extensive experiments demonstrate that our method accurately produces judgments conditioned on the input and template, and that later judgments can condition on earlier ones, enabling dependency-aware analysis. Furthermore, YOFO also benefits from post-hoc Chain-of-Thought (CoT), which encourages the model to reason implicitly before judging.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) [bai2025qwen2, wang2024qwen2, bai2023qwenvl, lu2025ovis2, lu2024ovis, hong2025worldsense, ma2024visual, ma2024multi, xue2023clip] have significantly extended the capabilities of Large Language Models (LLMs) [team2024qwen2, yang2025qwen3, ahmed2025qwen, bai2023qwen, liu2024deepseek, guo2025deepseek, li2025adaptive] by enabling them to perceive, reason over, and generate responses grounded in multimodal inputs. However, the dense visual token sequences required for such alignment incur prohibitive computational costs. Recent approaches address this by either compressing tokens while preserving fidelity or rethinking how visual information is integrated. For instance, CrossLMM [yan2025crosslmm] employs dual cross-attention mechanisms to drastically reduce token count with minimal performance degradation. Similarly, MM-GEM [ma2024multi] uses a PoolAggregator to support both generative and embedding tasks. These advances demonstrate that strategic token reduction can maintain multimodal performance at lower cost.
@@ 2.2 Rerankers
Reranking serves as a critical refinement stage in modern retrieval systems, aiming to re-order an initial candidate set by modeling fine-grained query–document interactions. Early neural rerankers employed representation-based architectures, calculating embeddings independently and then similarity, such as DSSM [huang2013learning, gao2014modeling], Convolutional DSSM [shen2014learning]) and Long Short-Term Memory (LSTM) variants [hochreiter1997long, palangi2016deep, wan2016deep, cohen2016end]. But this kind of models are limited by the fixed-size embedding, and thus were soon superseded by interaction-based cross-encoders, most notably BERT-based models, that concatenate query and document as input and predict relevance via a classification or regression head over the [CLS] token. [nogueira2019multi] first introduced this method and proved it effective. These models achieve strong performance but require pairwise inference, limiting scalability, and is constrained by fixed representational dimention and content length.   With the rise of LLMs, researchers have repurposed powerful encoder-decoder (e.g., T5 [raffel2020exploring]) or decoder-only (e.g., LLaMA [touvron2023llama]) backbones for reranking. A dominant paradigm [zhuang2023rankt5, ma2024fine] fine-tunes LLMs to regress continuous relevance scores or classify relevance levels, leveraging their rich contextual understanding and long-context capabilities for improved semantic matching. Jina-Reranker-M0 also follows this paradigm but leverages Qwen2-VL [wang2024qwen2], enabling support for multimodal inputs.   More recently, generative reranking has emerged as a paradigm shift: instead of scoring candidates individually, the model directly generates the ranked sequence—such as a list of document indices or IDs—conditioned on the full candidate set. Methods like [yoon2024listt5, yang2025rank] process all candidates jointly and autoregressively output the optimal ranking. This generative formulation naturally integrates reranking into end-to-end AI systems and shows particular promise in multimodal and agent-driven retrieval scenarios.
@@ 3.1 Problem formulation
As illustrated in Sec. 1, models akin to Jina-Reranker-M0 [jina2025blog] struggle to perform accurate cross-modal retrieval. One might attribute this to visual information degradation caused by the image encoder and the connector projection, which impairs fine-grained alignment. However, we empirically find that these models also fail to retrieve text accurately from natural-language queries. As shown in Tab. 1(a), Jina-Reranker-M0 assigns a higher score to winter trousers simply because they are gray, failing to recognize that the query explicitly requires summer-appropriate clothing, while “gray” is only a conditional preference. After further simplifying the prompt, Tab. 1(b) shows that Jina-Reranker-M0 fails to handle negation reliably, appearing to measure lexical overlap rather than semantic similarity on slightly more complex inputs. Verifying each requirement in the query individually would likely yield better results. To this end, we formally define the compositional judging problem.
@@ 4 Experiments
In this section, we present extensive experiments and ablation studies to validate our approach. We first compare YOFO with Jina-Reranker-M0 [jina2025blog], a representative state-of-the-art (SOTA) multimodal reranker, on the test set derived from LAION-RVS-Fashion [lepage2023lrvs]. We then assess YOFO’s dependency-aware judgement, the effectiveness of post-hoc CoT, and the contribution of individual architectural components.
@@ 4.1 Impementation Details
We build on Qwen2-VL-2B-Instruct [wang2024qwen2] and Qwen3-VL-2B-Instruct [qwen3vl2025blog], freezing the vision encoder during training. We fine-tune the LLM using LoRA [hu2022lora]. To conveniently locate the unknown token, we add four special tokens to the Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct processor: <|auth_start|>, <|auth_end|>, <|reason_start|>, and <|reason_end|>. The first two mark the start and end of the answer span, and the last two mark the boundaries of the reason span. Consequently, these special tokens enable straightforward label masking, allowing supervision to be applied exclusively to answers and reasons.
@@ 4.2 Data Construction
Training set. To construct the training dataset, we use the SA-1B dataset [kirillov2023segment], introduced alongside the Segment Anything Model. It contains approximately 11M diverse, high-resolution images sourced from a privacy-preserving data engine and annotated with 1.1B high-quality segmentation masks. Since we do not require the masks, as shown in Fig. 4(a), we use only the images, prompting an MLLM to analyze them, propose properties that hold or do not hold for the corresponding images, and provide reasons. The detailed prompt is shown in Fig. 5(a). We randomly sample 1.2M images to ensure diversity and generalizability. For each image $\mathcal{I}_{i}$, the final sample is represented as $(\mathcal{I}_{i},(\mathcal{P}_{ij},\mathcal{A}_{ij},\mathcal{R}_{ij})_{j=1}^{N})$, denoting the image, the proposed properties, the binary answers(“yes” or “no”), and the reasons. To maintain a balanced and randomized mix of satisfied and unsatisfied properties, we prompt the MLLM to produce approximately equal numbers of each and shuffle them during training. Additionally, although we instruct the model to generate exactly ten properties per sample in the training set, the number of properties per training sample varies, typically remaining close to ten.. To evaluate the model and mitigate overfitting, we split the dataset at a ratio of $6:1$ into training and validation sets.
Test set. We use the LRVS-Fashion dataset [lepage2023lrvs], a large-scale public benchmark for referred visual search (RVS) in the fashion domain, to construct a test set. It comprises 272K fashion products and 842K images extracted from real-world catalogs. To evaluate generalization in Image-Text reranking, as depicted as Fig. 4(b), we construct image pairs and queries; the model then judges which of the two images better matches the corresponding query. Specifically, we randomly sample image pairs $(\mathcal{I}_{i1},\mathcal{I}_{i2})$ from the same category and, using an MLLM, generate a natural-language query $\mathcal{Q}_{i}$ that mimics how a customer would express their preferences. The MLLM then produces the answer $\mathcal{A}_{i}$ indicating whether the first image $\mathcal{I}_{i1}$ better matches $\mathcal{Q}_{i}$, the set of query requirements $(\mathcal{R}_{ij})_{j=1}^{M_{i}}$ (where $M_{i}$ is the number of requirements), and a valid python expression $\mathcal{E}_{i}$ used to compute the final recommendation score based on our model’s predictions. The detailed prompt is demonstrated in Fig. 5(b).
@@ 4.4 Comparison on LAION-RVS-Fashion
In this section, we present an analysis of YOFO against state-of-the-art multimodal rerankers [jina2025blog].
@@
YOFO also delivers substantially higher throughput than generative rerankers. This efficiency derives from YOFO’s single-forward-pass design: unlike generative or multi-stage rerankers [ma2024fine, nogueira2019multi] that must process each requirement sequentially or in multiple rounds, YOFO produces all judgments in a single forward pass. Consequently, its inference latency increases only mildly with the number of requirements compared with generative or multi-stage rerankers, making it especially well suited to real-time, large-scale recommendation scenarios where both responsiveness and expressiveness are critical.
@@ 4.6 Ablation Study
Effectiveness of YOFO training. We fine-tune Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct models on the training set using LoRA [hu2022lora] for one epoch and evaluate it on the validation set. As shown in Tab. 4, after training, Property-wise Accuracy improves by over 20%, reaching above 90%, while Sample-wise Accuracy increases by approximately 40%. These ablation results demonstrate the effectiveness of our training paradigm. Moreover, the YOFO model trained with Qwen3-VL-2B-Instruct performs slightly better than its counterpart trained with Qwen2-VL-2B-Instruct. The Sample-wise Accuracy remains moderate because the training and validation sets contain, on average, nearly ten properties per sample, making it challenging to judge all properties correctly.